OpenStudy (anonymous):

If the point (2, -2) is on the terminal side of the angle θ, find each of sin(θ),cos(θ),tan(θ) si) = tan(θ) = @Mathematics

OpenStudy (anonymous):

sin(θ) = cos(θ) = tan(θ) =

OpenStudy (anonymous):

sin(θ)=y/r cos(θ)=x/r tan(θ)=y/x You can also draw a triangle  and use triangle trigonometry